OpenStudy (anonymous):

the slope of a line is -13/3. The point (-3, 7) is on the line. Determine the equation of this line

OpenStudy (anonymous):

Hey aroub :)

OpenStudy (aroub):

m=-13/2 (-3,7) y=mx+b 7=-13/2(-3)+b 7=39/2+b -39/2+14/2=b -25/2=b y=-13/3x-25/2 hey mathquiz!! :D

OpenStudy (anonymous):

tryna help me finish this submission theres like 5 more questions

OpenStudy (aroub):

YES SURE I LOVE SOLVING ABT SLOPES PROBS.! SO YES :)

OpenStudy (anonymous):

The slope of a line is 7/10. The point (-10, -9) is on the line. Determine the equation of this line

OpenStudy (aroub):

OKAY sry this time but u need to know how to slove it ill help u!! :)

OpenStudy (anonymous):

okay step by step aha

OpenStudy (aroub):

okay

OpenStudy (aroub):

thsi is so slow wanna go to twiddla??

OpenStudy (anonymous):

what is that??

OpenStudy (anonymous):

do you have aim?

OpenStudy (aroub):

http://www.twiddla.com/570310